Find the zeros of the equation. x^2-2x-15=y. I can't find anything in my book about finding zeros.


Found 2 solutions by nerdybill, mananth:
Answer by nerdybill(7384)

When they talk about finding "zeros", they are generally referring to x-intercepts. So, set y to zero and solve for x:
x^2-2x-15=y
x^2-2x-15=0
(x-5)(x+3) = 0
x = {-3, 5}
.
So, your zeros are at:
(-3,0) and (5,0)

Answer by mananth(16946)

Find the zeros of the equation. x^2-2x-15=y. I can't find anything in my book about finding zeros.
finding roots,finding zeros,finding factors of polynomials. -- All are same
So you know what it is now.
x^2-5x+3x-15=0
x(x-5)+3(x-5)=0
(x+3)(x-5)=0
(x+3)=0
OR
(x-5)=0
x= -3 OR x=5
...
m.ananth@hotmail.ca